Every student in a discrete math class is either a computer science or a mathematics major or is a joint major in these two subjects. How many students are in the class if there are 30 computer science majors(including joint majors), 24 math majors(including joint majors) and 9 joint majors?
Answer by checkley77(12844)

(30+24)-9=54-9=45 TOTAL STUDENTS.